If Kenya has a population of 20,000,000 people and doubling time of 16 years and if the growth continues at the same rate, find the population in 22 years.
Answer by checkley71(8403)

AN INCREASE OF 20,000,000 IN 16 YEARS IS PROPORTIONAL TO X GROWTH IN 22 YEARS.
20,000,000/16=X/22 CROSS MULTIPLY.
16X=22*20,000,000
16X=440,000,000
X=27,500,000 INCEREASE IN 22 YEARS CAN BE EXPECTED.
20,000,000+27,500,000=47,500,000 POPULATION IN 22 YEARS.